Operations Report – January 16th

Crew 219 Operations Report 16-JAN-2020

Sol: 11
Name of person filing report: Alejandro Perez

Non-nominal systems: N/A

Generator: Run
Hours run: 14
From what time last night: 17:30
To what time this morning: 07:30
List of any additional daytime hours when the generator was run: N/A

Solar- SOC% (Before the generator is run at night): 80%

09:10 = 100%; 10:00 = 99%;
11:00 = 97%;
13:00 = 88%;
14:00 = 85%;
15:00 = 82%;
16:30 = 80%;
18:00 = 100%

Diesel Reading – 55%
Station Propane Reading- 32%
Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – 50
Water Meter (units)- 0147767,3
Water (static tank) (gal) – 286
Static to loft Pump used –Yes
Water in Green Hab (gal): 149.03
Water in Science Dome (gal): 0

Toilet Tank Emptied –Yes

Deimos rover used: Still in the Workshop

Hours:
Beginning Charge:
End Charge:
Currently Charging:

Sojourner rover used: Assigned to Director
Hours:
Beginning Charge:
End Charge:
Currently Charging:

Spirit Rover used- Yes
Starting Hours: 126.1, 126.2
Beginning Charge: 100%, 100%
Ending Hours: 126.2, 126.4
Ending Charge: 51%, 76%

Opportunity Rover used: Still in the workshop
Hours:
Beginning Charge:
Ending Charge:
Currently Charging:

Curiosity Rover used: Yes
Starting Hours: 132.0, 132.1
Beginning Charge: 100%, 100%
Ending Hours: 132.1, 132.4
Ending Charge: 51%, 78%

Notes on Rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No
Reason for use:
Oil Added? No
# hours the ATV’s were used today:
Notes on ATVs:

Habcar used and why, where?
CrewCar used and why, where? In the shop.

General Notes and Comments:

Summary of the internet: Nominal

Summary of Suit and Radios: Suit #7 began a rapid power drop halfway through EVA 17.

Summary of Hab: Nominal

Summary of Science Dome operations: Average daytime temperature is 7°C. The average night time temperature is 2°C.

Summary of Ram operations: Nominal

Summary of any observatory issues: Nominal

Summary of Health and safety issues: Nominal

Questions, concerns, and requests to Mission Support:

EVA Report – January 16th

Crew 219 EVA Report 16-01-2020

EVA #17

Author: Dave Masaitis

Purpose of EVA: Sample return

Start time: 0908 hrs
End time: 1013 hrs

Narrative: EVA #17 completed a five-minute decompression by 0908 hours, mounted Spirit and Curiosity, and departed the Hab campus after reporting hours and SOC. The rovers proceeded northbound on Cow Dung Road until achieving a positive ID of Streambed Connector. The EVA team parked rovers and dismounted, proceeding east on foot along Streambed Connector. The streambed was full of patchy snow and ice, so the team moved cautiously along their directed route.

The route added approximately thirty minutes to planned transit time, so the team had to move swiftly to deposit samples once at their destination. Once complete, they discussed their route back and opted to try and retrace their incoming route as closely as possible. Only one team member fell through the ice, but no injuries were sustained, so they continued their movement to the rovers. The team proceeded southbound on Cow Dung Road to the Hab, re-entering the airlock at 1013 hours.

Destination: West of Candor Chasma
Coordinates: 0520000/4251500

Participants: D. Masaitis (HSO), N. Hadland (GHO), A. Elnajdi (GEO)

Roads and routes per MDRS Map: North on Cow Dung Road to Streambed Connector, East through Streambed Connector

Mode of travel: Rovers (Spirit and Curiosity) and Walking

EVA #18

Author: Cynthia Montanez

Purpose of EVA: Sample Return

Start Time: 1407 hrs
End Time: 1539 hrs

Narrative: EVA #18 is the last EVA crew 219 was able to do before completing their Mars analog mission. At 1407 hours, Cynthia (acting as CO) and Keith (acting as XO) lead their team 1 km south of Kissing Camel on Curiosity as Dave and Alejandro followed on Spirit.

When reaching their destination, Dave directed his crew to the specific locations each sample was taken from so that they could return their samples. Once completed, the EVA crew made their way back to the rovers and proceeded safely back to the Hab at 1539 hours.

Destination: 1 km S of Kissing Camel
Coordinates: 0518500/4248300

Participants: Cynthia Montanez (CO), Keith Crisman (XO), Dave Masaitis (HSO), Alejandro Perez (ENG)

Roads and Routes per MDRS Map: South on Cow Dung Road (0110), approximately 1 km south of Kissing Camel to return samples.

Mode of travel: Rovers (Spirit and Curiosity) and walking.

Sol Summary – January 16th

Crew 219 Sol Summary Report 16-01-2020

Sol: 11
Summary Title: If You Give a Martian a Cookie
Author’s name: Cynthia Montanez

Mission Status: Nominal

Sol Activity Summary: The morning started with crew 219 sipping on coffee and indulging in a light breakfast before they partook in their nominal tasks. To train C. Montanez and K. Crisman for their next field season, C. Montanez took the position as CO and K. Crisman took the position as XO for the day. Between 0908 to 1018, EVA #17 entered sim for sample return and came back to a delicious dish of pork sausage meatballs and spaghetti made by H. Blackburn.

Once everyone finished eating, each person started to work on finishing up their projects, post-mission report edits, and/or cleaning up parts of the Hab. Between 1402 to 1534, EVA #18 entered sim for the final sample return and came back to freshly made chocolate chip oatmeal cookies, brownies, and chai tea made by H. Blackburn. The aroma of the desserts made the Hab smell like grandma’s house, as the taste brought the crew back to Earth.

Finally, the day ended with H. Blackburn cooking coconut curry and crew positions being switched back. Overall, today was a great day on Mars.

Look Ahead Plan: Cleaning the Hab and its exterior facilities in preparation for the next crew

Weather: Chilly and sunny all-day

Crew Physical Status: Nominal

EVA:

EVA #17 – Sample Return W of Candor Chasma (0520000/4251500)
EVA #18 – Sample Return 1 km S of Kissing Camel (0518500/4248300)

Reports to be filed:

Sol Summary
EVA Reports #17 and #18
Operations Report
GreenHab Report
Research Report
Astronomy Report
Journalist’s Report
Photos (Including Photo of the Day)

Support Requested: None

Operations Report – January 15th

  

Operations Report

Crew 219 Operations Report 15-JAN-2020

Sol: 10

Name of person filing report: Alejandro Perez

Non-nominal systems: N/A

Generator: Run

Hours run: 14

From what time last night: 16:00

To what time this morning: 06:00

List of any additional daytime hours when the generator was run: N/A

Solar- SOC% (Before the generator is run at night): 76%

10:00 = 100%; 11:30 = 98%; 12:30 = 95%; 14:00 = 90%; 16:10 = 83%; 18:00 = 78%

Diesel Reading – 60%

Station Propane Reading- 35%

Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – 50

Water Meter (units)- 0147714,1

Water (static tank) (gal) – 328

Static to loft Pump used –Yes

Water in Green Hab (gal): 161.21

Water in Science Dome (gal): 0

Toilet Tank Emptied –No

Diemos rover used: Still in the Work shop

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Sojourner rover used: Assigned to Director

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Spirit Rover used- Yes

Starting Hours: 125.6

Beginning Charge: 100%

Ending Hours: 126.1

Ending Charge: 51%

Opportunity Rover used: Still in workshop

Hours:

Beginning Charge:

Ending Charge:

Currently Charging:

Curiosity Rover used: Yes

Starting Hours: 131.6

Beginning Charge: 100%

Ending Hours: 132.0

Ending Charge: 51%

Notes on Rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use:

Oil Added? No

# hours the ATV’s were used today:

Notes on ATVs:

Habcar used and why, where? Used by Outpost for the Post Office and to
acquire an oil pump for the generator.

CrewCar used and why, where? In the shop.

General Notes and Comments:

Summary of the internet: Nominal

Summary of Suit and Radios: All suits have been tested for battery
performance. All suit batteries output higher voltage than the
replacement batteries. Suits 1 and 6 have slightly lower air flow on
the right fan, however, this did not cause any major problems during
EVA.

Summary of Hab: Nominal

Summary of Science Dome operations: Temperature averaging 7°C during
the day and 0°C at night.

Summary of Ram operations: Nominal

Summary of any observatory issues: Nominal

Summary of Health and safety issues: Nominal

Questions, concerns, and requests to Mission Support:

EVA Report – January 15th

Crew 219 EVA Report 15-01-2020
EVA #15
Author: Dave Masaitis
Purpose of EVA: Sampling of geologic evidence of life forms
Start time: 1001 hrs
End time: 1147 hrs
Narrative: EVA #15 exited the Hab campus northbound on foot towards
North Ridge at 1001 hrs. According to the briefed plan, D. Masaitis
moved approximately 100m in advance of the rest of the EVA element to
gain visual identification of an unimproved route and begin recording
coordinates for future use. Upon reaching the top of the route, D.
Masaitis halted to await the rest of the team, and upon their arrival,
EVA #15 began identifying sites of interest. The EVA team sampled
along the northern and western limits of Hab Ridge first, and then
moved south and east as they cataloged samples. Upon reaching the
eastern limit overlooking the Hab, the team conversed and concluded
that they had appropriately screened the area, and decided to proceed
carefully down the eastern slope of Hab Ridge to arrive at the Hab by
1147 hours.
Destination: Hab Ridge
Coordinates: 0517700/4251600
Participants: D. Masaitis (CO), K. Crisman (HSO), H. Blackburn (LSO),
C. Montanez (GHO)
Roads and routes per MDRS Map: N towards North Ridge, WSW up an
unimproved route ( coordinates listed below) adjacent to North Ridge,
S to Hab Ridge, and E back to the Hab.

Coordinates for unimproved route, from NE to SW (Safe for future passage):

0518298/4251366
0518260/4251374
0518211/4251369
0518164/4251376
0518116/4251334
0518076/4251317
0518018/4251294
0517961/4251275

Crew 219 EVA Report 15 JAN 2020

EVA#: 16

Author: Nathan Hadland

Purpose of EVA: The primary purpose of this EVA was sampling for our
geology project and astrobiology projects. Post-EVA, we collected data
for our biometrics project.

Start time: 1313 hrs

End time: 1542 hrs

Narrative: At 1313 hours, EVA #16 initiated a 5-minute decompression
and moved to Curiosity and Spirit. After reporting SOC and hours, we
progressed south along Cow Dung Road until we reached the intersection
of Peiades Rd. We moved east on foot towards the North Pinto Hills.
Along the way, we took 5 regolith samples and recorded coordinates.
The crew then continued northeast to the North Pinto Hills region. The
crew then took additional biological samples along the way. The crew
then turned around and headed southwest back to the rovers. Curiosity
and Spirit safely arrived back at the Hab at 1537 hours and
successfully depressurized in the airlock at 1542 hours.

Destination: North Pinto Hills

Coordinates: 0521000/4248800

Participants: Nathan Hadland (XO), Alejandro Perez (ENG), Robinson
Raphael (ASTRO), Abdul Elnagdi (GEO)

Roads and routes per MDRS Map: South on Cow Dung Road (0110), parking
at the intersection of Peiades Rd and walking northeast towards North
Pinto Hills.

GreenHab Report – January 15th


Crew 219 GreenHab Report 15-JAN-20

Crew GreenHab Officer: Cynthia Montanez

Environmental control: Heating.

Shade cloth (40% and 30%) on.

Average temperature: 24.3 °C

2:34 P.M.

Floor Unit: 18 °C

Mid-height unit: 26 °C

4:18 P.M.

Floor Unit: 18 °C

Mid-height unit: 24 °C

6:06 P.M.

Floor Unit: 14 °C

Mid-height unit: 24 °C

Hours of supplemental light: Light system 7:00PM-12:00AM.

Daily water usage for crops: 9.18 Gal.

Daily water usage for research and/or other purposes: N/A

Water in Blue Tank – 152.03 Gal.

Time(s) of watering for crops:

2:34 P.M.

SEASONAL PLANTING HISTORY:

Change to crops: N/A

OTHER NOTES:

1st sprouts: One of the recently planted snow peas on the middle rack
is starting to sprout.

Harvest: At 4:35 P.M., 1 gram of grapefruit mint and 7 grams of chives
were harvested.

Other: N/A

Sol Summary – January 15th

Crew 219 Sol Summary Report 15-01-2020

Sol: 10

Summary Title: What day is it, camel?

Author’s name: Dave Masaitis

Mission Status: Nominal

Sol Activity Summary: The crew filtered out casually for coffee and a
light breakfast, before EVA #15 moved downstairs to suit up. The team
moved out towards North Ridge, documenting their route up to Hab
Ridge, so that future teams can use the same route. They found a
multitude of viable samples and returned to the Hab in time to have a
quick lunch and help EVA #16 get out on time. While EVA #16 moved out
for sampling, those at the campus processed samples and started
drafting language for their respective sections of the end of mission
summary. H. Blackburn prepared cheddar biscuits and hot chocolate for
the crew to enjoy as EVA #16 returned, and crew members got to work
preparing their evening reports. H. Blackburn prepared stir-fry in the
wok for dinner, and after the Comms window, the crew looks forward to
turning command over to C. Montanez and K. Crisman for a Thursday of
on-the-job command experience so they will have a taste of what is to
come next year. D. Masaitis will spend the day assuming the role of
Health and Safety Officer, and N. Hadland will take over as GreenHab
Officer. Their Command-exchange will end after tomorrow’s Comms
window, but will better prepare them to understand the roles they will
be returning to fulfill next year.

Look Ahead Plan: Two sample return EVA’s, materials consolidation,
revisions of mission summary

Anomalies in work: None

Weather: Chilly and sunny all day

Crew Physical Status: Nominal

EVA:

EVA #15 – Sample Collection on Hab Ridge (0517700/4251600)

EVA #16 – Sample Collection around North Pinto Hills (0521000/4248800)

Reports to be filed:

Sol Summary

EVA Report #15 & 16

EVA Request #17 & 18

Operations Report

GreenHab Report

Research Report

Journalist’s Report

Photos (Including Photo of the Day)

Support Requested:

None

Journalist Report – January 15th


Crew 219 Journalist Report

Author: Nathan Hadland, Crew 219 Executive Officer

Sol 10

It’s Wednesday My Crews

The ship is sailing on smooth seas, as they say. The EVAs today went
smoothly, all of our projects progressed forward, and we had a few
good laughs along the way. Our crew is beginning the process of
tightening up our science, consolidating samples, and writing up final
reports. We have collected a great deal of interesting data, explored
fascinating regions and landscapes, and have had a lot of fun in the
process. I am proud with what our crew has accomplished so far and am
excited to see what our post-mission analysis brings.

We had two EVAs today, both of which accomplished a great deal. The
first EVA mounted Hab Ridge along Sagan Street to collect lichen and
gryphaea fossils. Unfortunately for me, our mission timeframe landed
in the middle of graduate school application season. While my fellow
crewmembers were on EVA, I was laboring away at those application
essays. Oh well. The price I pay for trying to do too many things at
once.

We had a tight timeline today, so as soon as the first EVA arrived
back at the Hab and post-EVA biometrics data was collected, we began
preparations for our second EVA of the day. I led the expedition to
the North Pinto Hills to collect additional geologic samples as well
as lichen. The sun was especially potent today, in both its brightness
and heat. As soon as we began trudging through the snow and mud, we
started to feel our solar companion’s effects. We climbed a short way
and Abdul (GEO) identified a few sites for sample collection.
Afterwards, the team collectively realized the immense beauty of the
surrounding landscape and took the opportunity to capture a few
photographs.

The EVA team then headed further northeast to search for samples of
astrobiological interest. I’ve got to applaud Alejandro (ENG) and his
impeccable eye for finding lichen hidden amongst the rocks. After
climbing another hill looking for sampling sites, we decided to head
back to the rovers. Along the way, we stumbled across some incredible
geologic features and amazing views of the surrounding landscape. I
took the opportunity to teach some of my crewmembers orienteering and
navigation skills.

Upon arrival back at the Hab, we were greeted with hot chocolate and
warm homemade biscuits prepared by Hannah (LSO). Oh man, they were
heavenly. Tonight, we will be feasting on stir fry and chai tea.
Tomorrow, Dave (CO) and I will be training Cynthia (GHO) and Keith
(HSO) on commanding officer duties for when they return to MDRS next
season. Overall, I am extremely pleased and proud of our crew’s
science and am very happy with the group of talented individuals I
have chosen to surround myself with.

Operations Report – January 15th

  

Operations Report

Crew 219 Operations Report 14-JAN-2020

Sol: 10

Name of person filing report: Alejandro Perez

Non-nominal systems: N/A

Generator: Run

Hours run: 14

From what time last night: 16:00

To what time this morning: 06:00

List of any additional daytime hours when the generator was run: N/A

Solar- SOC% (Before the generator is run at night): 76%

10:00 = 100%; 11:30 = 98%; 12:30 = 95%; 14:00 = 90%; 16:10 = 83%; 18:00 = 78%

Diesel Reading - 60%

Station Propane Reading- 35%

Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – 50

Water Meter (units)- 0147714,1

Water (static tank) (gal) – 328

Static to loft Pump used –Yes

Water in Green Hab (gal): 161.21

Water in Science Dome (gal): 0

Toilet Tank Emptied –No

Diemos rover used: Still in the Work shop

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Sojourner rover used: Assigned to Director

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Spirit Rover used- Yes

Starting Hours: 125.6

Beginning Charge: 100%

Ending Hours: 126.1

Ending Charge: 51%

Opportunity Rover used: Still in workshop

Hours:

Beginning Charge:

Ending Charge:

Currently Charging:

Curiosity Rover used: Yes

Starting Hours: 131.6

Beginning Charge: 100%

Ending Hours: 132.0

Ending Charge: 51%

Notes on Rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use:

Oil Added? No

# hours the ATV’s were used today:

Notes on ATVs:

Habcar used and why, where? Used by Outpost for the Post Office and to
acquire an oil pump for the generator.

CrewCar used and why, where? In the shop.

General Notes and Comments:

Summary of the internet: Nominal

Summary of Suit and Radios: All suits have been tested for battery
performance. All suit batteries output higher voltage than the
replacement batteries. Suits 1 and 6 have slightly lower air flow on
the right fan, however, this did not cause any major problems during
EVA.

Summary of Hab: Nominal

Summary of Science Dome operations: Temperature averaging 7°C during
the day and 0°C at night.

Summary of Ram operations: Nominal

Summary of any observatory issues: Nominal

Summary of Health and safety issues: Nominal

Questions, concerns, and requests to Mission Support: